OpenStudy (anonymous):

Sharon kicks a ball from the ground into the air with an upward velocity of 64 feet per second. Write a function that models the height h, in feet, of the ball at time t, in seconds. When will the ball reach the ground again? A. 1 second after the ball is thrown B. 2 seconds after the ball is thrown C. 3 seconds after the ball is thrown D. 4 seconds after the ball is thrown Projectile motion: h(t)= -16t^2+V_0t+h_0 Please help!! :(

OpenStudy (anonymous):

-16t^2+64t+0?

OpenStudy (anonymous):

hello :)

OpenStudy (welshfella):

use the equation of motion for a projectile moving un der gravity h = ut + (1/2)gt^2 where u = initial velocity, h = height, t = time and g = acceleration due to gravvity here u = 64 fts-1 and g = -32 ft s-2

OpenStudy (danjs):

1D path, only force is gravity. initial conditions at h=0 and V = 64

OpenStudy (welshfella):

h = 64t - 16t^2

OpenStudy (welshfella):

the ball will reach the ground when h = 0 so you need to solve 64t - 16t^2 = 0

OpenStudy (welshfella):

one solution will be t = 0 which corresponds to when the ball is first kicked

OpenStudy (anonymous):

And the other is 4??

OpenStudy (welshfella):

64t - 16t^2 = 0 16t(4 - t) = 0 4-t = 0 t = 4 Yes